With current technology, suppose a firm is producing 500 shirts daily. Also assume that the least-cost combination of resources in producing those shirts is 10 units of labor, 8 units of land, 3 units of capital, and 4 units of entrepreneurial ability, selling at prices of $50, $40, $60, and $30, respectively. If the firm can sell these 500 shirts at $2 per unit, what is its total revenue? Its total cost? Its profit or loss? Will it continue to produce shirts? If this firm’s situation is typical for the other makers of shirts, will resources flow toward or away from this apparel?
Answer #1

The least cost combination of producing the shirts is 10 units of labour, 8 units of land, 3 units of capital and 4 units of entrepreneurial ability. The factors of production are sold at a price of $50, $40, $60 and $30 respectively. Taking into consideration, the factors of production mentioned here are at a per unit basis.

The total revenue after selling the shirts = $2*500 = $1000.

The total cost of making the shirts = 10*50+8*40+3*60+4*30= 500+320+180+120 = $1120.

The total cost here is more than its total revenue. So, the firm will run in a loss.

It will not continue to produce shirts as long as the price if shirts per unit for not rise which could have overcome the cost and incurred profit. As long as the price if each shirt remains to be $2 per unit, and the cost of production remains totbe the same, firm will not further continue to produce shirts as this will continuously incur loss. Hence, firm will not continue to produce shirts further.

Resources will not flow toward this apparel, as the firm is making losses, hence other firms letting selling the factors of production to the shirt firm will tend to shift their allocation of resources to other firms which are profit maximizers. As these firms already know the shirt firm is incurring losses, and the t-shirt firms could close down any time soon. So, the firms producing the factors of production will not risk themselves and will shift away their resources to allocate them to other firms.